# A Perfectly Truthful Calibration Measure

### v1

@@ Abstract
Calibration requires that predictions are conditionally unbiased and, therefore, reliably interpretable as probabilities. Calibration measures quantify how far a predictor is from perfect calibration. As introduced by Haghtalab et al. (2024), a calibration measure is truthful if it is minimized in expectation when a predictor outputs the ground-truth probabilities. Although predicting the true probabilities guarantees perfect calibration, in reality, when calibration is evaluated on a finite sample, predicting the truth is not guaranteed to minimize any known calibration measure. All known calibration measures incentivize predictors to lie in order to appear more calibrated on a finite sample. Such lack of truthfulness motivated Haghtalab et al. (2024) and Qiao and Zhao (2025) to construct approximately truthful calibration measures in the sequential prediction setting, but no perfectly truthful calibration measure was known to exist even in the more basic batch setting.
We design a perfectly truthful calibration measure in the batch setting: averaged two-bin calibration error (ATB). In addition to being truthful, ATB is sound, complete, continuous, and quadratically related to two existing calibration measures: the smooth calibration error smCal and the (lower) distance to calibration $\underline{\textsc{distCal}}$. The simplicity in our definition of ATB makes it efficient and straightforward to compute. ATB allows faster estimation algorithms with significantly easier implementations than smCal and $\underline{\textsc{distCal}}$, achieving improved running time and simplicity for the calibration testing problem studied by Hu et al. (2024). We also introduce a general recipe for constructing truthful measures based on the variance additivity of independent random variables, which proves the truthfulness of ATB as a special case and allows us to construct other truthful calibration measures such as quantile-binned $\ell_{2}$-ECE.
@@ 1 Introduction
Probabilistic forecasting has become increasingly important in modern AI-assisted decision-making. Unlike deterministic classification, probabilistic forecasts provide uncertainty quantification, allowing assessment of risks. One desired property of probabilistic predictions is calibration, which requires predictions to be conditionally unbiased and, therefore, reliably interpretable as probabilities. For example, neural networks for tumor diagnosis are trained to output a prediction $r\in[0,1]$, ideally interpretable as the probability of a binary state $y\in\{0,1\}$: the tumor segment being malignant or not. The neural network is calibrated if, conditioned on the output being, say $40\%$, the probability that the tumor is malignant is also $40\%$.
@@
There is a recent line of work studying the truthfulness of calibration measures (Haghtalab et al., 2024; Qiao and Zhao, 2025) in the sequential prediction setting. We focus on the more basic batch setting, where a calibration measure is truthful if it is minimized, in expectation over a random sample, when a predictor outputs the ground truth probabilities (we term this predictor the truthful predictor). Ideally, to assess calibration, a calibration measure should rank a calibrated predictor with the lowest expected error. This would imply that the calibration measure is truthful: the truthful predictor is always calibrated and should achieve minimum calibration error. However, no known calibration measure is truthful. It is true that assuming knowledge of the full joint distribution of the prediction and the state, all known calibration measures reach the minimum 0 if and only if a predictor is perfectly calibrated, but in practice, calibration error is evaluated on random samples rather than the unobserved underlying distribution. On a random sample of arbitrary size, for any known calibration measure, predicting the true probabilities does not guarantee minimum expected error, and a lower expected error does not guarantee that the predictor is more calibrated.
@@ 1.1 Non-truthfulness of Known Calibration Measures
We illustrate the non-truthfulness of ECE in Example 1.1 below, with a discussion afterwards showing that the same non-truthfulness issue occurs in a broad family of calibration measures. We consider the batch setting: a sample of $T$ individuals whose binary states $\bm{y}=(y_{1},\ldots,y_{T})\in\{0,1\}^{T}$ are independently drawn from the Bernoulli distributions with means $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ (denoted by $\bm{y}\sim\bm{p}$). We say $\bm{p}=(p_{1},\ldots,p_{T})$ are the ground-truth probabilities. In the example below, a predictor strictly benefits from reporting less informative, pooled predictions than the ground-truth probabilities.
Example 1.1 (ECE and binned ECE are not truthful, c.f. Qiao and Valiant, 2021).
@@
Suppose the ground-truth probabilities are $\bm{p}=(p_{1},\dots,p_{T})$ where each $p_{t}$ is distributed independently and uniformly over $[1/3,2/3]$. An uninformative predictor that always predicts $r_{1}=\cdots=r_{T}=0.5$ achieves an expected empirical $\textsc{ECE}=O\left(\sqrt{1/T}\right)$, the sampling error. However, a truthful predictor who reports $r_{t}=p_{t}$ results in a higher empirical $\textsc{ECE}\geq 1/3$. This is because the predictions $r_{1},\ldots,r_{T}\in[1/3,2/3]$ are almost surely distinct, so the empirical conditional expectation $\widehat{r}_{t}:=\operatorname{\mathbb{E}}_{(r,y)\sim\mathrm{Unif}((r_{t},y_{t})_{t\in[T]})}[y|r=r_{t}]$ is simply $y_{t}\in\{0,1\}$, giving $|r_{t}-\widehat{r}_{t}|=|r_{t}-y_{t}|\geq 1/3$.
The same observation holds for binned ECE as well. Suppose the prediction space of $[0,1]$ is discretized into $k$ equal-length intervals (i.e., bins). When the true probabilities are predicted, on average, each interval has roughly $\frac{T}{k}$ predictions, inducing a sampling error of roughly $\sqrt{k/T}$, higher than the binned ECE $O\left(\sqrt{1/T}\right)$ achieved by the uninformative predictor $r_{1}=\cdots=r_{T}=0.5$.
@@
Example 1.1 shows that the truthful predictor $\bm{r}=\bm{p}$ does not achieve a minimum ECE or binned ECE on a sample of size $T$. Even worse, there are miscalibrated predictors (e.g. the predictor that always predicts $0.5+\epsilon$ for a small $\epsilon>0$) achieving smaller ECE and binned ECE than the calibrated truthful predictor. Thus ECE and binned ECE do not rank predictors correctly based on how calibrated they are.
@@
The observation in Example 1.1 that predicting the truth incurs higher (i.e. worse) calibration error than (strategically) assigning the same prediction (i.e. the overall average) to every individual is, unfortunately, very general. It happens not just for ECE and its variants [^1] Variants of ECE include $\ell_{\alpha}$-ECE, where we replace the absolute bias |rr^||r-$\widehat${r}| with $|r-\widehat{r}|^{\alpha}$ for an arbitrary $\alpha\geq 1$, as well as binned versions of $\ell_{\alpha}$-ECE., but also for continuous calibration measures [^2] These are calibration measures that are continuous as a function of the predictions. Note that ECE and binned ECE are not continuous. such as the smooth calibration error (Kakade and Foster, 2008), the distance to calibration (Błasiok et al., 2023) and its variants, etc, irrespective of the sample size $T$ used for estimating the calibration error. Moreover, it happens consistently accross every realization of the states $\bm{y}$, not just in expectation. Specifically, for every realization of states $\bm{y}\in\{0,1\}^{T}$ and every prediction sequence $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$, all these calibration measures Cal satisfy
@@
where $\bar{\bm{r}}=(\bar{r},\ldots,\bar{r})\in[0,1]^{T}$ is the constant predictor that always predicts $\bar{r}:=\frac{1}{T}\sum_{t=1}^{T}r_{t}$. This means that regardless of the (unknown) realized states $\bm{y}$, the constant predictor $\bar{\bm{r}}$ consistently achieves smaller or equal calibration error than $\bm{r}$, even when $\bm{r}$ is the true underlying probabilities $\bm{p}$. We formally prove this observation in Theorem A.2.
It is very easy to find (many) examples of $\bm{y}$ where the inequality (1) becomes strict, in which case reporting $\bar{\bm{r}}$ instead of the truth $\bm{r}=\bm{p}$ is strictly better. In particular, we can find many examples of $\bm{y}$ where $\textsc{Cal}(\bm{r},\bm{y})>0$ and $\textsc{Cal}(\bar{\bm{r}},\bm{y})=0$. This is because for all of the calibration measures Cal mentioned above (including the continuous ones), we always have $\textsc{Cal}(\bm{r},\bm{y})>0$ as long as $r_{1},\ldots,r_{T}$ are distinct values in, say, $[1/3,2/3]$, and we have $\textsc{Cal}(\bar{\bm{r}},\bm{y})=0$ as long as the average outcome $\frac{1}{T}\sum_{t=1}^{T}y_{t}$ is equal to $\bar{r}$.
1.2 Problem Statement
@@
Measuring and optimizing for calibration non-truthfully only makes the predictions less trustworthy, going in the very opposite direction of the goal of calibration. Let $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ be the ground-truth probabilities. We say a calibration measure Cal is truthful if outputting $\bm{p}$ minimizes the expected empirical calibration error, i.e., the minimizer to $\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{r},\bm{y})]$ is $\bm{r}=\bm{p}$, for every $\bm{p}\in[0,1]^{T}$. From a machine learning perspective, a truthful measure helps identify the Bayesian optimal predictor (Gneiting and Raftery, 2007) because it correctly ranks ground-truth predictions with the lowest expected error. From a game-theoretic perspective, a truthful measure incentivizes an optimizing predictor to output their true beliefs, where we view $\bm{p}$ as the predictor’s subjective belief about the probabilities, which might differ from the true probabilities. For the tumor risk prediction task, if assessed by a non-truthful calibration measure, a doctor is incentivized to report a prediction different from their true probabilistic assessment of the tumor risk, to make the predictions look more calibrated. In this case, an incentivized misreport can hardly be trusted.
@@
In the sequential prediction setting, where the states $y_{1},\ldots,y_{T}$ are revealed sequentially after each prediction $r_{t}$ is made, Haghtalab et al. (2024) and Qiao and Zhao (2025) design approximately truthful calibration errors, where predicting the true probabilities approximately minimizes the expected error. The simpler batch setting, where all predictions are made before the states are revealed, is left unexplored. We observe that, in the batch setting, some existing measures, such as the smooth calibration error (Kakade and Foster, 2008; Błasiok et al., 2023) and the calibration measures proposed by Haghtalab et al. (2024) and Qiao and Zhao (2025), are approximately truthful. Yet, no known calibration measure is perfectly truthful.
Our goal is to design perfectly truthful calibration measures in the batch setting. Such measures correctly rank the truthful predictor with the lowest expected error and incentivize a predictor to output their true beliefs about the state distributions.
@@
Applying a truthful error metric alone does not solve our problem. A calibration measure should, at a minimum, be able to distinguish calibrated predictors from miscalibrated ones given a sufficiently large sample. That is, it should satisfy basic properties, termed completeness and soundness (Definitions 2.8 and 2.9), requiring vanishing error when a predictor is calibrated (completeness) and non-vanishing error when a predictor is miscalibrated (soundness) as the sample size increases. Observed by Haghtalab et al. (2024), some error metrics, such as the well-known squared error $\frac{1}{T}\sum_{t}(r_{t}-y_{t})^{2}$, are truthful but far from being a complete and sound calibration measure.
As our main result, we construct a perfectly truthful, complete and sound calibration measure in the batch setting. Notably, our calibration measure provides a stronger rank-preserving guarantee than perfect truthfulness: the expected error over samples is the same as the idealized calibration error on the unknown true underlying distribution, up to a constant term independent of the predictor. In particular, it correctly ranks all calibrated predictors (not just the truthful one) with the same lowest expected error.
Moreover, we show that our calibration measure maintains and even improves other desiderata in the recent research on constructing good calibration measures beyond completeness, soundness, and truthfulness: consistency (Błasiok et al., 2023), continuity, low sample complexity, and high computational efficiency. Our calibration measure satisfies consistency in Błasiok et al., 2023, requiring it to be polynomially related to the Distance to Calibration, defined as the $\ell_{1}$ distance to the closest calibrated predictor. Our measure is continuous as a function of the predictions. It has the same low sample complexity $O(1/\varepsilon^{2})$ for estimation as the smooth calibration error. It has a very simple definition, making it more efficient to compute than existing continuous measures such as the smooth calibration error and the Distance to Calibration, achieving improved running time and simplicity for the calibraition testing problem studied by Hu et al. (2024). In summary, our construction achieves the best of all worlds while strictly improving many desired properties (e.g. truthfulness, computational efficiency, and simplicity).
1.3 Our Contributions
@@
As our main contribution, we construct a perfectly truthful, complete and sound calibration measure: Averaged Two-Bin Calibration Error ($\mathrm{ATB}$). In addition, ATB satisfies all the other desired properties discussed earlier: continuity, low sample complexity, high computational efficiency, and the rank-preserving property. Moreover, it is quadratically related to two existing calibration measures: the smooth calibration error smCal (Kakade and Foster, 2008) and the (lower) Distance to Calibration $\underline{\textsc{distCal}}$ (Błasiok et al., 2023). This relationship implies that ATB is a consistent calibration measure in the framework of Błasiok et al. (2023).
As a bonus benefit of its simplicity and computational efficiency, ATB gives a faster algorithm for testing calibration compared to the previous algorithms of Hu et al. (2024). The $\ell_{1}$ version of ATB gives a simple constant-factor approximation to smCal and $\underline{\textsc{distCal}}$ while being more efficiently computable.
@@
We also introduce a general framework for constructing truthful error metrics that encompasses ATB as a special case and allows us to construct other truthful calibration measures such as quantile-binned $\ell_{2}$-ECE.
@@
Our construction of $\mathrm{ATB}$ is very simple. We first divide the predictions into just two bins $[0,q)$ and $[q,1]$, with the bin boundary $q$ chosen uniformly at random from $[0,1]$. Within each bin, we compute the squared error between the sum of the predictions and the sum of the states, divided by $T^{2}$. After summing up the errors in the two bins, we define $\mathrm{ATB}$ to be the expectation over the random choice of the bin boundary $q$:
@@
We show that ATB is truthful and continuous (in fact, Lipschitz continuous) as a function of the predictions $\bm{r}$:
Theorem 1.3 (Informal, Theorems 4.3 and 4.4).
$\mathrm{ATB}$ is a truthful and continuous calibration error. Here, truthful means predicting the ground truth $\bm{p}$ achieves lower or equal expected error than any other report vector $\bm{r}$:
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{p},\bm{y})]\leq\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r},\bm{y})]\quad\text{for every $\bm{p},\bm{r}\in[0,1]^{T}$}.$
Continuity means that for every two prediction sequences $\bm{r}_{1},\bm{r}_{2}\in[0,1]^{T}$,
@@
We show that ATB is quadratically related to two known continuous, complete and sound calibration measures: the smooth calibration error smCal (Kakade and Foster, 2008) and the lower distance to calibration $\underline{\textsc{distCal}}$ (Błasiok et al., 2023). We include formal definitions of smCal and $\underline{\textsc{distCal}}$ later in Definitions 2.5 and 2.4. Roughly speaking, the lower distance to calibration $\underline{\textsc{distCal}}(\bm{r},\bm{y})$ is the $\ell_{1}$ earthmover distance from the reported predictions $\bm{r}$ to the closest perfectly calibrated predictions. The smooth calibration error smCal is the following supremum over $1$-Lipschitz functions $w:[0,1]\to[-1,1]$:
$\textsc{smCal}(\bm{r},\bm{y})=\sup_{\text{$1$-Lipschitz }w:[0,1]\to[-1,1]}\frac{1}{T}\sum_{t\in[T]}w(r_{t})\cdot(r_{t}-y_{t}).$
Błasiok et al. (2023) showed that smCal and $\underline{\textsc{distCal}}$ are both continuous, complete and sound calibration measures, and they are constant-factor approximations of each other.
Theorem 1.4 (see Corollary 5.2).
For any sequences $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$ and $\bm{y}=(y_{1},\ldots,y_{T})\in\{0,1\}^{T}$,
$\frac{2}{9}\textsc{smCal}(\bm{r},\bm{y})^{2}$ $\leq\mathrm{ATB}(\bm{r},\bm{y})\leq 6\textsc{smCal}(\bm{r},\bm{y}),\text{\quad and\quad}$
$\frac{1}{18}\underline{\textsc{distCal}}(\bm{r},\bm{y})^{2}$ $\leq\mathrm{ATB}(\bm{r},\bm{y})\leq 3\underline{\textsc{distCal}}(\bm{r},\bm{y}).$
Theorem 1.4 implies that ATB is complete, sound, and is a consistent calibration measure in the framework of Błasiok et al. (2023).
We also show that ATB has desirable sample complexity and supreme computational complexity:
Theorem 1.5 (Informal, see Theorem 4.5 and Section 4.4).
ATB can be estimated within $\varepsilon$ error using $T=O(\varepsilon^{-2})$ examples, and can be computed in $O(T\log T)$ time.
The $O(T\log T)$-time algorithm for computing ATB follows directly from the simplicity of its definition. We need $O(T\log T)$ time to sort the predictions once, after which it only takes $O(T)$ time (see Section 4.4). For comparison, the best-known algorithms for computing smCal and $\underline{\textsc{distCal}}$ take $O(T\log^{2}T)$ and $O(T^{2}\log T)$ time, respectively (Hu et al., 2024).
More Than Truthfulness: the Rank-Preserving Property of ATB.
We show that ATB has the following stronger property that implies truthfulness. It allows us to compare predictors based on calibration on the sampled states and produce the same ranking as if calibration is evaluated on the unobserved underlying distribution:
Theorem 1.6 (Informal, see Theorem 4.3).
Let $\bm{p}\in[0,1]^{T}$ denote the true probabilities. For two prediction sequences $\bm{r}_{1},\bm{r}_{2}\in[0,1]^{T}$, we have
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r}_{1},\bm{y})]-\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r}_{2},\bm{y})]=\mathrm{ATB}(\bm{r}_{1},\bm{p})-\mathrm{ATB}(\bm{r}_{2},\bm{p}).$
Here, on the left-hand side, we compare the expected empirical error $\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r},\bm{y})]$ on the sampled states $\bm{y}$, whereas on the right-hand side, we compare the error $\mathrm{ATB}(\bm{r},\bm{p})$ w.r.t. the unobserved true probabilities $\bm{p}$. The theorem guarantees consistent ordering between these two ways of evaluating calibration. The truthfulness property in Theorem 1.3 then follows immediately as a consequence of Theorem 1.6 (and the basic fact that $\mathrm{ATB}(\bm{p},\bm{p})=0$):
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r},\bm{y})]-\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{p},\bm{y})]=\mathrm{ATB}(\bm{r},\bm{p})-\mathrm{ATB}(\bm{p},\bm{p})\geq 0.$
In particular, Theorem 1.6 implies that any calibrated predictions $\bm{r}$ w.r.t. the true probabilities $\bm{p}$ (i.e. predictions with zero Distance to Calibration $\underline{\textsc{distCal}}(\bm{r},\bm{p})=0$, or equivalently $\mathrm{ATB}(\bm{r},\bm{p})=0$ due to their quadratic relationship) achieve the same minimum expected empirical error $\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r},\bm{y})]$ as predicting the truth, because
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r},\bm{y})]-\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{p},\bm{y})]=\mathrm{ATB}(\bm{r},\bm{p})-\mathrm{ATB}(\bm{p},\bm{p})=0-0=0.$
The converse is also true and similarly implied by Theorem 1.6.
@@
We define UBSEs as binning-based calibration measures where the binning scheme can be randomized and can depend arbitrarily on the predictions $\bm{r}$. The key to its truthfulness lies in how the biases in each bin are combined to calculate the final error. We compute the squared biases in each bin similarly to the standard $\ell_{2}$ ECE, but there is a subtle but crucial difference in how these squared biases are weighted and combined. We discuss this family in more detail in Section 1.4.1 and Section 3. We prove that each UBSE has the following key property of error decomposition, which implies truthfulness as well as the rank-preserving property in Theorem 1.6:

### v2

@@ Abstract
Calibration requires that predictions are conditionally unbiased and, therefore, reliably interpretable as probabilities. A calibration measure quantifies how far a predictor is from perfect calibration. As introduced by Haghtalab et al. (2024), a calibration measure is truthful if it is minimized in expectation when a predictor outputs the ground-truth probabilities. Predicting the true probabilities guarantees perfect calibration, but in reality, when calibration is evaluated on a random sample, all known calibration measures incentivize predictors to lie in order to appear more calibrated. Such lack of truthfulness motivated Haghtalab et al. (2024) and Qiao and Zhao (2025) to construct approximately truthful calibration measures in the sequential prediction setting, but no perfectly truthful calibration measure was known to exist even in the more basic batch setting.
We design a simple, perfectly and strictly truthful, sound and complete calibration measure in the batch setting: averaged two-bin calibration error (ATB). ATB is quadratically related to two existing calibration measures: the smooth calibration error smCal and the lower distance to calibration $\underline{\textsc{distCal}}$. The simplicity in our definition of ATB makes it efficient and straightforward to compute, allowing us to give the first linear-time calibration testing algorithm, improving a result of Hu et al. (2024). We also introduce a general recipe for constructing truthful measures based on the variance additivity of independent random variables, which proves the truthfulness of ATB as a special case and allows us to construct other truthful calibration measures such as quantile-binned $\ell_{2}$-ECE.
@@ 1 Introduction
Probabilistic forecasting has become increasingly important in modern AI-assisted decision-making. Unlike deterministic classification, probabilistic forecasts provide uncertainty quantification, allowing assessment of risks. One desired property of probabilistic predictions is calibration, which requires predictions to be conditionally unbiased and, therefore, reliably interpretable as probabilities. For example, neural networks for tumor diagnosis are trained to output a prediction $r\in[0,1]$, ideally interpretable as the expectation of a binary state $y\in\{0,1\}$: the tumor segment being malignant or not. The neural network is calibrated if, conditioned on the output $r$ being, say $40\%$, the probability that the tumor is malignant $\Pr[y=1|r=40\%]$ is also $40\%$.
@@
A recent line of work studies the truthfulness of calibration measures (Haghtalab et al., 2024; Qiao and Zhao, 2025). An error measure is truthful if it incentivizes a predictor to output the truth, i.e., the expected error is minimized when the predictor reports the true probabilities. However, no known calibration measure is truthful. Even a miscalibrated predictor can have lower expected error than the truthful predictor when evaluated by known calibration measures. We explain this non-truthfulness in Section 1.1.
@@ 1.1 Non-truthfulness of Known Calibration Measures
For existing non-truthful calibration measures, there exists obvious and uninformative prediction strategies that induce a lower calibration error than truth-telling. Concretely, the “obvious” strategy is to ignore all features and always predict the base rate, i.e., output the same constant for every input. In this section, we explain why this obvious strategy has a strictly lower error in expectation, and a lower loss even for every realization of samples. Following this observation, it is well-known that calibration is a poor measure of accuracy (DeGroot and Fienberg, 1983).
@@
Example 1.1 illustrates the non-truthfulness of ECE. We consider the batch setting: a sample of $T$ individuals whose binary states $\bm{y}=(y_{1},\ldots,y_{T})\in\{0,1\}^{T}$ are independently drawn from the Bernoulli distributions with means $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ (denoted by $\bm{y}\sim\bm{p}$). We say $\bm{p}=(p_{1},\ldots,p_{T})$ are the ground-truth probabilities. In the example below, a predictor strictly benefits from reporting the base rate of ground truths.
Example 1.1 (ECE is not truthful, c.f. Qiao and Valiant, 2021).
@@
Suppose the ground-truth probabilities are $\bm{p}=(p_{1},\dots,p_{T})$ where each $p_{t}$ is distributed independently and uniformly from $[1/3,2/3]$. An uninformative predictor that always predicts $r_{1}=\cdots=r_{T}=0.5$ achieves an expected empirical $\textsc{ECE}=O\left(\sqrt{1/T}\right)$, the sampling error. However, a truthful predictor who reports $r_{t}=p_{t}$ results in a higher empirical $\textsc{ECE}\geq 1/3$. This is because the predictions $r_{1},\ldots,r_{T}\in[1/3,2/3]$ are almost surely distinct, so the empirical conditional expectation $\widehat{r}_{t}:=\operatorname{\mathbb{E}}_{(r,y)\sim\mathrm{Unif}((r_{t},y_{t})_{t\in[T]})}[y|r=r_{t}]$ is simply $y_{t}\in\{0,1\}$, giving $|r_{t}-\widehat{r}_{t}|=|r_{t}-y_{t}|\geq 1/3$.
@@
Example 1.1 shows that the obvious uninformative prediction achieves a lower ECE. Even worse, there are miscalibrated predictors (e.g. the predictor that always predicts $0.5+\epsilon$ for a small $\epsilon>0$) achieving smaller ECE than the calibrated truthful predictor. Thus ECE do not rank predictors correctly based on how calibrated they are.
@@
Predicting the uninformative base rate incurs a lower calibration error for all known calibration measures. It happens not just for ECE in Example 1.1 and its variants [^1] Variants of ECE include $\ell_{\alpha}$-ECE, where we replace the absolute bias |rr^||r-$\widehat${r}| with $|r-\widehat{r}|^{\alpha}$ for an arbitrary $\alpha\geq 1$, as well as binned versions of $\ell_{\alpha}$-ECE., but also for continuous calibration measures [^2] These are calibration measures that are continuous as a function of the predictions. Note that ECE and binned ECE are not continuous. such as the smooth calibration error (Kakade and Foster, 2008), the distance to calibration (Błasiok et al., 2023) and its variants, etc, irrespective of the sample size $T$. Moreover, it happens consistently accross every realization of the states $\bm{y}$, not just in expectation. Specifically, for every realization of $\bm{y}\in\{0,1\}^{T}$ and every prediction sequence $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$, all these calibration measures Cal satisfy
@@
where $\bar{\bm{r}}=(\bar{r},\ldots,\bar{r})$ is the constant predictor that always predicts the average $\bar{r}:=\frac{1}{T}\sum_{t=1}^{T}r_{t}$. This obvious and uninformative strategy always achieves (weakly) lower calibration error. For many realizations of states $\bm{y}$, the error is strictly lower. We formally prove this observation in Theorem A.2.
1.2 Our Goal: Perfectly Truthful Calibration Measures
@@
Measuring and optimizing for calibration non-truthfully only makes the predictions less trustworthy, going in the very opposite direction of the goal of calibration. Recall the definition of truthfulness: a calibration measure Cal is truthful if for every ground-truth probabilities $\bm{p}\in[0,1]^{T}$, the expected empirical calibration error $\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{r},\bm{y})]$ of predictions $\bm{r}\in[0,1]^{T}$ on a random sample $\bm{y}\sim\bm{p}$ is minimized when $\bm{r}=\bm{p}$. From a machine learning perspective, a truthful measure helps identify the Bayes optimal predictor (Gneiting, 2011) because it correctly ranks ground-truth predictions with the lowest expected error. From a game-theoretic perspective, a truthful measure incentivizes an optimizing predictor to output their true beliefs, where we view $\bm{p}$ as the predictor’s subjective belief about the probabilities, which might differ from the true probabilities. For the tumor risk prediction task, if assessed by a non-truthful calibration measure, a doctor is incentivized to report a prediction different from their true probabilistic assessment of the tumor risk, to make the predictions look more calibrated. Such an incentivized misreport can hardly be trusted.
@@
We focus on the perfect truthfulness in the batch setting. Previous work (Haghtalab et al., 2024; Qiao and Zhao, 2025) design approximately truthful calibration measures in the sequential prediction setting, where the states $y_{1},\ldots,y_{T}$ are revealed sequentially after each prediction $r_{t}$ is made. We observe that, in the simpler batch setting, some existing measures are approximately truthful, such as the smooth calibration error (Kakade and Foster, 2008; Błasiok et al., 2023) and the calibration measures proposed by Haghtalab et al. (2024) and Qiao and Zhao (2025). Yet, no known calibration measure is perfectly truthful.
@@
The two minimum requirements of a calibration measure are completeness and soundness. A calibration measure should be able to distinguish calibrated predictors from miscalibrated ones given a sufficiently large sample. Completeness requires vanishing error when a predictor is calibrated (Definition 2.8), and soundness requires non-vanishing error when a predictor is miscalibrated as the sample size increases (Definition 2.9). Haghtalab et al. (2024) point out that some error metrics, such as the well-known squared error $\frac{1}{T}\sum_{t}(r_{t}-y_{t})^{2}$, are truthful but far from being a complete and sound calibration measure. The squared error of a calibrated predictor may not vanish as the sample size $T$ increases.
Our main result shows that truthfulness can be achieved via surprisingly simple constructions in the batch setting, while preserving the completeness and soundness of existing calibration measures.
1.3 A Roadmap of Our Contributions
@@
As our main contribution, we construct a perfectly truthful, complete and sound calibration measure: Averaged Two-Bin Calibration Error ($\mathrm{ATB}$). In addition, ATB maintains and even significantly improves other desired properties considered in the literature: continuity and consistency (Błasiok et al., 2023), low sample complexity, and high computational efficiency.
@@
We propose a general framework for constructing truthful error metrics, where ATB is a special case. This general framework allows us to construct other truthful calibration measures while preserving existing properties. For example, we construct the truthful quantile-binned $\ell_{2}$-ECE with surprisingly simple changes to binned ECE.
We identify key properties that must be satisfied by every truthful measure: it must preserve the Blackwell ordering of calibrated predictors and must (weakly) decrease in expectation when we recalibrate a miscalibrated predictor. Since our $\mathrm{ATB}$ and quantile-binned $\ell_{2}$-ECE are truthful, they inherit these properties and can be used as meaningful and robust evaluation metrics for predictors in practice.
As bonus side products of our simple truthful construction, we give the first linear-time algorithm for the calibration testing problem studied by Hu et al. (2024) as well as a simple, faster constant-factor approximation to the smooth calibration error and the Lower Distance to Calibration.
@@
Section 4 introduces ATB, constructed with two bins and a randomized binning boundary:
  • The prediction space of $[0,1]$ is divided into two bins $[0,q)$ and $[q,1]$, with the bin boundary $q$ chosen uniformly at random from $[0,1]$.
  • Within each bin, we compute the squared error between the sum of the predictions and the sum of the states.
  • After summing up the errors in the two bins and dividing the result by $T^{2}$, we define $\mathrm{ATB}$ to be the expectation over the random choice of the bin boundary $q$.
@@
We prove the following properties of ATB.
  • Perfect truthfulness (Theorem 4.8). We additionally show that ATB is strictly truthful, meaning that truth-telling is the only minimizer of the expected error, in the ex-ante stage where a predictor is evaluated on a random sample of $T$ i.i.d. individuals.
  • Lipschitz continuity in predictions. For every two prediction sequences $\bm{r}_{1},\bm{r}_{2}\in[0,1]^{T}$,
@@
  • Completeness and soundness. The completeness and soundness of ATB follow from the quadratic relationship to two known continuous, complete and sound calibration measures: the smooth calibration error smCal (Kakade and Foster, 2008) and the lower distance to calibration $\underline{\textsc{distCal}}$ (Błasiok et al., 2023). See Theorem 4.3.
  • Sample and computational complexity. ATB can be estimated within $\varepsilon$ error using $T=O(\varepsilon^{-2})$ examples in $O(T)$ time. For exact computation of ATB, Theorem 4.12 shows an $O(T\log T)$-time algorithm on $T$ examples. For approximate computation, Theorem 4.12 shows an $O(T+1/\varepsilon)$-time algorithm for approximating up to additive error $\varepsilon>0$. For $\varepsilon\approx 1/\sqrt{T}$, the running time is $O(T+1/\varepsilon)=O(T)$.
    These algorithms are simple and easy to implement, following the simplicity of the definition of ATB. For exact computation, our algorithm sorts the predictions in $O(T\log T)$ time, and does the rest of the computation in $O(T)$ time. For $\varepsilon$-additive approximation, the sorting step can be implemented in $O(T+1/\varepsilon)$ time by discretizing the predictions before sorting. For comparison, the best-known algorithms for computing smCal and $\underline{\textsc{distCal}}$ are much more complicated and take $O(T\log^{2}T)$ and $O(T^{2}\log T)$ time, respectively, even when $O(1/\sqrt{T})$ additive error is allowed (Hu et al., 2024).
@@
We define UBSEs as binning-based calibration measures where the binning scheme can be randomized and can depend arbitrarily on the predictions $\bm{r}$. The key to its truthfulness lies in how the biases in each bin are combined to calculate the final error. We compute the squared biases in each bin similarly to the standard $\ell_{2}$ ECE, but there is a subtle but crucial difference in how these squared biases are weighted and combined. We discuss this family in more detail in Section 1.4.1 and Section 3. We prove that each UBSE has the following key property of error decomposition, which implies truthfulness.